A cob Douglas production function for a firm is given as Q=4L ½K½. The firm has also established that wage rate and interest paid on capital are $3 and $5 respectively for a production period. The firm intents to spend $200 million for the period on production cost. Compute the levels of capital and labor that will maximize output. What is the maximum output?                      (10 Marks)


1
Expert's answer
2021-10-31T18:31:54-0400

Form a Lagragian equation

"Q=4""\left(L^{\smash{0.5}}\right)\left(K^{\smash{0.5}}\right)" subject to "wL+rK=C"

"L=4" "\left(L^{\smash{0.5}}\right)\left(K^{\smash{0.5}}\right)" "-\u03bb(wL+rK-C)"

"\u03b4L\/\u03b4L=2""\left(L^{\smash{-0.5}}\right)\left(K^{\smash{0.5}}\right)""-\u03bbw=0..........(i)"

"\u03b4L\/\u03b4K=2" "\left(L^{\smash{0.5}}\right)\left(K^{\smash{-0.5}}\right)" "5-\u03bbr=0...........(ii)"

"\u03b4L\/\u03b4\u03bb=wL+rK-C=0.................(iii)"

Divide equation (i) and (ii)

"K\/L=w\/r" and thus  "K=wL\/r" and "L=rK\/w"

Replacing the two equation on equation "(iii)"

"w(kr\/w)+rK=C" thus "K*=C\/2r"

"\left(K^{\smash{*}}\right)=200/\left( 2^{\smash{*}}\right) 5=20"

"wL+r(wL\/r)=C"

"wL+wL=C"

"\left(L^{\smash{*}}\right)=C/2w"

"\left(L^{\smash{*}}\right)=200/\left(2^{\smash{*}}\right)3 = 33.33"

Hence the optimum output is;

"Q=4" "\left(33.33^{\smash{0.5}}\right)\left(20^{\smash{0.5}}\right)"

"Q=4(4.58)(4.5)"

"Q=104.40" units - Maximum Output.